# All-fiber microendoscopic polarization sensing at single-photon level aided by deep-learning

### v2

@@ 1.3 Numerical setup
$F=\text{Tr}\left[\sqrt{\sqrt{\rho}\cdot\sigma\cdot\sqrt{\rho}}\right]^{2}\in\left[0,1\right],$ (S3)
@@
2 1951 USAF birefringent test
@@
We performed an additional analysis beyond those presented in the main text. Specifically, we conducted a scan of a 1951 USAF birefringent resolution test with the developed polarization sensor and compared it with a polarization measurement based on rotating wave plates. Fig. S2 shows a polarization image obtained with the standard method (a) alongside the scan using our all-fiber sensor (b). Both panels visualize false-colored RGB images using Bloch parameters $\left(B_{1},B_{2},B_{3}\right)$ with $R=\frac{B_{1}+1}{2},G=\frac{B_{2}+1}{2},$ and $B=\frac{B_{3}+1}{2}$ corresponding to red, green, and blue channel intensities. By averaging over the background and polarization-modulating regions, we estimated a fidelity value of approximately 0.986 between the two polarization measurement approaches. This comparison unequivocally demonstrated the ability of our sensor to perform highly accurate polarization measurements.
Additionally, the image shown in panel (b) allows for the evaluation of the spatial resolution of our sensor. To this end, we fitted horizontally and vertically oriented edges at several positions. Characterized as the 20% to 80% width, we estimated the resolution to be approximately $(6\pm 5)\leavevmode\nobreak\ \mu$m. Besides the distance from a specimen to the fiber tip, the resolution is dictated by the few-mode fiber diameter and can be modified by employing different fiber types. Moreover, tapering the fiber tip could improve the resolution of our sensor up to the near-field scanning optical microscopy level.
@@
Figure S2: A comparison of polarization measurement on a region of the 1951 USAF birefringent resolution test. (a) A polarization image obtained using standard imaging based on rotating wave plates. (b) An all-fiber sensor scan of the same region. Both results are visualized as false-colored RGB images using Bloch parameters.
3 Dense connective tissue
In the main text, we demonstrated our sensor by performing a polarization scan of dense connective tissue. For comparative purposes, Fig. S3 includes an additional panel depicting a polarization image of the same specimen obtained using a stand-alone polarization microscope. Its experimental setup comprises a monochrome camera (ImagingSource DMK 23U274) and a pair of crossed polarizers. The typical inner structure of the dense connective tissue is evident in both polarization images, confirming that our sensor indeed performs polarization characterization of this sample. Moreover, the false-color image provided by our sensor contains complete information about the polarization states.
Figure S3: Visualization of dense connective tissue: (left) an intensity image, (middle) a scan using the all-fiber polarization sensor, (right) an image using a stand-alone polarization microscope. The first two panels reproduce results from the main text. The additional panel quantitatively verifies the tissue polarization structures via a common crossed-polarizers microscope.

### v3

@@ 4.3 Numerical setup
$F=\text{Tr}\left[\sqrt{\sqrt{\rho}\cdot\sigma\cdot\sqrt{\rho}}\right]^{2}\in\left[0,1\right],$ (3)
@@
4.4 1951 USAF birefringent test
@@
We performed an additional analysis beyond those presented in the main text. Specifically, we conducted a scan of a 1951 USAF birefringent resolution test with the developed polarization sensor and compared it with a polarization measurement based on rotating wave plates. Fig. 6 shows a polarization image obtained with the standard method (a) alongside the scan using our all-fiber sensor (b). Both panels visualize false-colored RGB images using Bloch parameters $\left(B_{1},B_{2},B_{3}\right)$ with $R=\frac{B_{1}+1}{2},G=\frac{B_{2}+1}{2},$ and $B=\frac{B_{3}+1}{2}$ corresponding to red, green, and blue channel intensities. By averaging over the background and polarization-modulating regions, we estimated a fidelity value of approximately 0.987 between the two polarization measurement approaches. This comparison unequivocally demonstrated the ability of our sensor to perform highly accurate polarization measurements.
@@
Figure 6: A comparison of polarization measurement on a region of the 1951 USAF birefringent resolution test. (a) A polarization image obtained using standard imaging based on rotating wave plates. (b) An all-fiber sensor scan of the same region. Both results are visualized as false-colored RGB images using Bloch parameters.
Additionally, the image shown in panel (b) allows for the evaluation of the spatial resolution of our sensor. To this end, we fitted horizontally and vertically oriented edges at several positions. Characterized as the 20% to 80% width, we estimated the resolution to be approximately $(6\pm 6)~{}\mu$m. Besides the distance from a specimen to the fiber tip, the resolution is dictated by the few-mode fiber diameter and can be modified by employing different fiber types. Moreover, tapering the fiber tip could improve the resolution of our sensor up to the near-field scanning optical microscopy level.
Declarations
[…]
Availability of data and materials
The code and data that support the findings of this study are publicly available on GitHub: https://github.com/VasinkaD/Polarization-Deep-Sense.
[…]
Funding
This work was supported by the Czech Science Foundation (project 21-18545S), Ministry of Education, Youth, and Sports of the Czech Republic (project OP JAC CZ.02.01.01/00/23_021/0008790). MB and DV acknowledge the support by Palacký University (project IGA-PrF-2024-008).
[…]
Authors’ contributions
MB assembled the sensor and conducted the experiments. DV wrote the deep learning and data processing codes, and drafted manuscript. MJ initiated and supervised the project. All authors contributed in data interpretation and revising manuscript.
[…]
Acknowledgements
We acknowledge the use of cluster computing resources provided by the Department of Optics, Palacký University Olomouc. We thank Jan Provazník for maintaining the cluster and providing support. We also thank Josef Hloušek for developing a 3D model shown in Fig. 1 and Jaromír Běhal for a fruitful discussion on birefringent biological samples.